# Experimental Ground-State Energy of a 125-Site Flat Kagome Antiferromagnet via Hamiltonian Engineering on Quantum Computer

### v2

@@ III.3 Physical Interpretation of J’ = 2
Although a full dimer loop-flip offers a precise route to ground-state accuracy, it becomes increasingly difficult to simulate such global flips as the Kagome lattice expands over multiple unit cells. This limitation arises due to the hardware constraints and non-local mapping of neighboring sites in the ansatz. Nevertheless, as demonstrated in Figure 1, setting $J^{\prime}=2$ on a single bond of a defect triangle within a 6-site subregion can sufficient to leave more potent signatures of the spin-liquid states. With $J^{\prime}=2$, the same ansatz circuit enables entanglement among qubits 2, 3, 4, and 5, effectively unfreezing the static dimer cover into a superposition state that better approximates the local physics of a spin liquid. This localized resonance significantly reduces energy estimation error, even with shallow circuits and limited entanglement connectivity.
@@ IV The Proposed VQE Algorithm
To address scalability and circuit depth limitations, we decompose the full 100-qubit VQE into two stages: local VQE and global VQE. In the local phase, we optimize the parameters of smaller ansatz circuits (15–19 qubits showsn in Figure-3) that correspond to segments of the Kagome lattice. Each subproblem is independently solved on a classical simulator using exact diagonalization or small-scale VQE, yielding accurate estimates of local energy contributions. These local solutions collectively define a baseline energy, which serves as an upper bound for the global VQE. The global phase, executed on a quantum processor, adjusts the interconnections between pre-optimized segments, allowing for long-range entanglement that can lower the total energy and approach the true ground-state of the full system.
@@ IV.1 Local VQE: Classical Computing
Figure 8 illustrates the qubit-to-site mapping and its decomposition into seven local VQE segments. Each segment covers a chain of linearly adjacent, corner-sharing triangles, and includes two defect triangles, with the exception of segment 7, which contains only one. This segmentation strategy reflects both the topological structure of the Kagome lattice and the constraints of the hardware-efficient ansatz, allowing localized parameter optimization within manageable sub-circuit sizes.
A key objective of the local VQE phase is to determine an appropriate enhancement of the exchange coupling $J^{\prime}$ on selected bonds of each defect triangle. For every lattice segment, the value of $J^{\prime}$ is individually calibrated to ensure that the computed local ground-state energy matches the exact ground-state energy obtained via exact diagonalization. As shown in Table 2, the calibrated values of $J^{\prime}$ remain remarkably consistent across all subregions, clustering around $J^{\prime}\approx 1.95$. This empirical regularity supports the scalability of our Hamiltonian engineering approach to larger systems.
At the conclusion of the local VQE phase, we obtain a baseline estimate of the ground-state energy gs_energy as $-161$. In the subsequent phase, the sub-circuits that contributed to this value are recombined into a complete ansatz for global VQE, with the goal of further lowering the energy.
@@ IV.2 Global VQE: Quantum Computing
Local VQE produces pre-optimized ansatz segments for each subregion of the Kagome lattice. To recover the full 103-qubit circuit, these segments are concatenated such that all original gate and qubit dependencies of the ansatz are preserved, effectively reconstructing the near original circuit. However, this naïve concatenation does not guarantee global optimality because the junction qubits—shared between two segments—inherit independently optimized $R_{y}(\theta)$ rotations from each side. To address the potential sub-optimality introduced at these junctions, we insert an additional trainable $R_{y}(\theta)$ gate on each junction qubit (see Fig. 3), and tune these parameters via a global VQE.
@@ IV.2.2 Objective Function
Concatenating seven local ansatz segments introduces six free parameters corresponding to the newly inserted $R_{y}(\theta)$ gates on the junction qubits (e.g., qubits 17, 33, 48, 63, 73, 92). The global VQE runs the reconstructed ansatz on IBM quantum processors, with the objective of minimizing $\text{Tr}(\rho H_{\text{pert}})$, where $H_{\text{pert}}$ denotes the engineered Hamiltonian derived from the KAFH model.
$H_{SEL}=J\sum_{(i,j)\in S}X_{i}X_{j}+Y_{i}Y_{j}+Z_{i}Z_{j}$
@@
Since the full Hamiltonian includes over 568 noisy Pauli terms, we adopt a noise-aware simplification by retaining only those terms involving the junction qubits and their immediate neighbors—this reduced Hamiltonian is denoted $H_{\text{SEL}}$. While this truncation reduces the expressiveness of the cost function, it significantly improves signal-to-noise ratio. Owing to the entanglement path through the 102 CNOT gates of the hardware-efficient ansatz, all qubits remain technically interdependent, and thus parameter updates at a single junction can, in principle, affect the global quantum state.
@@ IV.3 Noise Mitigation with Operator Decoherence Renormalization (ODR)
To implement this correction, we consider a parameter regime in which the VQE ansatz circuit becomes efficiently classically simulable—typically by reducing to a Clifford circuit. Importantly, the Clifford version of the ansatz arises solely due to specific parameter choices; the underlying circuit structure (i.e., the gates and their order) remains unchanged. We refer to this special configuration as the reference ansatz.
In our setting, the Clifford (reference) version of the ansatz prepares a static dimer cover state $\langle{\phi}$, making the expectation value
$\langle H\rangle=\langle\phi|H|\phi\rangle$
the ideal (noise-free) output of the reference circuit. When the same reference ansatz is executed on the quantum device, it produces a noisy (mixed) state $\sigma$, yielding the expectation
$\operatorname{Tr}(\sigma H_{\text{pert}}).$
In contrast, the actual ansatz used in the VQE algorithm typically operates in a non-Clifford regime, as it explores the full parameter space to minimize energy. Nevertheless, once the VQE optimization is complete, the ODR technique can be applied post hoc to mitigate the noise in the final energy estimate.
@@
$\frac{\text{noise-mitigated}\operatorname{Tr}(\rho H_{\text{pert}})}{\operatorname{Tr}(\rho H_{\text{pert}})}=\frac{\operatorname{Tr}(\sigma H)}{\langle{\phi}|H|\phi\rangle}$ (1)
@@ V Experimental Results
Table 5 presents the optimal parameter sets obtained from the global VQE runs on ibm_aachen, ibm_fez, and ibm_torino. These optimal parameters appear to be significantly influenced by processor-specific factors, including calibration quality and error rates. Nevertheless, there is no fundamental limitation preventing cross-testing of these parameter sets across different processors, which can help expand the search space for identifying the lowest gs_energy estimate. Therefore, treating Table 5 as a general candidate pool, we proceed to compute the noise-mitigated $\textnormal{Tr}(\rho H_{pert})$ by trial assignment of each parameter set to the blue-shaded gates in Figure-4, across all processors.
The values reported in Table 3 were obtained by selecting, for each processor, the parameter set from Table 5 that yields the minimum noise-mitigated $\textnormal{Tr}(\rho H_{pert})$. To ensure consistency, we aim to maximize the ratio $\frac{\textnormal{Tr}(\rho H_{pert})}{\textnormal{Tr}(\sigma H)}$ within the same Qiskit job. This strategy avoids device-level behavioral and calibration fluctuations that can occur when measurements are spread across multiple Qiskit estimator jobs. The numerical values in Table 3 are based on at least three independent repetitions of the same configuration (circuit, parameter set, and estimator settings). See Figure 4(d) for details on the estimator settings. Table 3 summarizes energy estimates from five different quantum processors, all falling within the range [–165.91, –172.32]. The lowest value corresponds to a per-site energy of approximately –0.415 for the spin-1/2 KAFH model.
For comparison, classical simulations using Qiskit’s Aer backend with the Matrix Product State (MPS) method and circuit cutting [36] yield estimates between –167 and –168, closely matching quantum results. This validates the accuracy of our noise-mitigated global VQE. To benchmark our results, we introduce heuristic energy bounds. The system contains 60 triangles, each contributing –3 in the ideal dimerized configuration, giving a theoretical lower bound of –180. However, defect triangles systematically raise the energy. For example, a single defect triangle on the smallest sub-lattice (two-unit cells as in Figure-7(d)) elevates gs_energy by 0.86 (Table-1(d)). With 13 such defects, and each contributing $\approx$+0.86  the effective upper bound becomes –168.82. This value aligns closely with our best estimates from both quantum and classical simulations, reinforcing the accuracy of our Hamiltonian-engineered VQE strategy.
Table 3: The unmitigated (Actual) and noise-mitigated (Noise mitigated) gs_energy using ODR, for five different ibm quantum processors
Processor Actual $\textnormal{Tr}(\rho H_{pert})$ Ref. ODR $\textnormal{Tr}(\sigma H)$ Noise mitigated $\textnormal{Tr}(\rho H_{pert})$
@@ V.2 Comparison of Classically-optimized and Quantum-optimized Parameters
To ensure a fair comparison, both parameter sets were independently evaluated on three quantum processors—ibm_torino, ibm_marrakesh, and ibm_fez, by plugging them into the full 103-qubit ansatz and measuring the ground-state energy without applying any noise mitigation (e.g., ODR). The unmitigated energy values obtained from $\mathbf{p}_{1}$ were $-121.47\pm 1.05$, $-105.58\pm 0.72$, and $-115.77\pm 0.46$, respectively. In contrast, the energy values using $\mathbf{p}_{2}$ were consistently lower: $-122.81\pm 0.53$, $-106.63\pm 0.33$, and $-117.32\pm 0.22$. Although the numerical differences are moderate, these results suggest that parameters learned from quantum VQE directly on hardware not only reflect the underlying circuit structure but also adaptively account for calibration specifics and coherent noise characteristics [32], potentially navigating toward lower-energy configurations. This finding supports the view that quantum-optimized parameters can outperform those obtained from classical approximations, particularly when hardware-specific factors significantly influence algorithm convergence.
Table 4: Comparison of unmitigated ground-state energies using parameters from classical VQE ($\mathbf{p}_{1}$) and quantum VQE ($\mathbf{p}_{2}$) on different IBM quantum processors.
Processor Energy with $\mathbf{p}_{1}$ (J) Energy with $\mathbf{p}_{2}$ (J) Average Energy Decrease (J)
@@
An important objective of this study is to demonstrate utility-grade quantum computation. We have already shown that the parameter set $\texttt{p}_{2}$, obtained via quantum execution, slightly outperforms the classical counterpart $\texttt{p}_{1}$ in one-to-one comparisons on specific processors. Figure-5 presents a more holistic, statistically informed comparison of global VQE parameter performance on quantum processors with that of $\texttt{p}_{1}$ on classical computer. It contrasts the quantum gs_energy estimates, averaged over all parameter sets listed in Table 5 and all five processors, with the classical baseline estimate using $\texttt{p}_{1}$ (i.e., $-167.95$).
@@
As quantum hardware continues to improve in qubit fidelity and coherence, quantum validation of classically approximated quantum algorithms opens an exciting avenue for future research. It provides a concrete benchmark for assessing the reliability of classical simulations, especially as quantum processors scale to larger qubit counts with increasing gate fidelity. Our work presents an initial step toward this promising direction and encourages further exploration of its potential.
@@ VI Benchmarking Against the Thermodynamic Ground State Energy
In this work, we present what is, to our knowledge, the first experimental attempt to estimate the ground-state energy of a 103-site KAFH lattice on superconducting quantum hardware. By employing a Hamiltonian engineering strategy, we modulate bond strengths to favor valence bond configurations and optimize energy estimation using a hardware-efficient VQE ansatz. On IBM quantum processors, our estimate for the ground-state energy per site of a 125-site flat KAFH lattice falls in the range $-0.405$ to $-0.415J$, depending on the specific device utilized. This result lies marginally above the thermodynamic-limit energy, which has been determined via large-scale DMRG and tensor network methods to lie between $-0.436$ and $-0.438J$ per site [13, 12].
@@
A substantial part of this discrepancy arises from finite-size and boundary effects, which are well known to raise energy estimates in systems with open boundary conditions. In our lattice geometry, 24 out of 125 sites lie on the boundary and possess only two nearest neighbors, as opposed to the four neighbors found in the bulk. This reduced local coordination diminishes quantum fluctuations and suppresses dimer resonance, leading to locally elevated bond energies and an increased average energy per site.
To account for these edge effects, we apply an open-boundary correction (OBC) model, following established procedures from prior finite-size studies [28, 14]. By decomposing the total energy into edge and bulk contributions, and conservatively assuming an average energy of $-0.33J$ for edge sites—consistent with under-coordinated site estimates from DMRG and series expansion—we estimate the corrected bulk energy density to be approximately $-0.429J$ per site. This corrected value falls within the expected thermodynamic range and supports the conclusion that our approach, combining Hamiltonian engineering with a resource-efficient VQE ansatz, effectively captures the essential physics of the KAFH model within current hardware constraints.

### v3

@@ III.3 Physical Interpretation of $J^{\prime}\approx 2$
Although a full dimer loop-flip offers a precise route to ground-state accuracy, it becomes increasingly difficult to simulate such global flips as the Kagome lattice expands over multiple unit cells. This limitation arises due to the hardware constraints and non-local mapping of neighboring sites in the ansatz. Nevertheless, as demonstrated in Figure-1, setting $J^{\prime}=2$ on a single bond of a defect triangle within a 6-site subregion can sufficient to leave more potent signatures of the spin-liquid states. With $J^{\prime}=2$, the same ansatz circuit enables entanglement among qubits 2, 3, 4, and 5, effectively unfreezing the static dimer cover into a superposition state that better approximates the local physics of a spin liquid. This localized resonance significantly reduces energy estimation error, even with shallow circuits and limited entanglement connectivity.
@@ IV The Proposed VQE Algorithm
To address scalability and circuit depth limitations, we decompose the full 103-qubit VQE into two stages: local VQE and global VQE. See Figure-3 for details. In the local phase, we optimize the parameters of smaller ansatz circuits (15–19 qubits) that correspond to subregions of the Kagome lattice displayed in Figure-8. Each VQE sub-problem is independently solved on a classical simulator yielding accurate estimates of local energy contributions and the relevant rotation gate $R_{y}(\texttt{c})$ parameter: c (unshaded gates in Figure-3). These local solutions collectively define a baseline energy, which serves as an upper bound for the global VQE. The global phase, executed on a quantum processor, adjusts the interconnections between pre-optimized segments, allowing for long-range entanglement that can lower the total energy and approach the true ground-state of the full system.
@@ IV.1 Local VQE: Classical Computing
Figure-8 illustrates the qubit-to-site mapping and its decomposition into seven local VQE segments. Each segment covers a chain of linearly adjacent, corner-sharing triangles, and includes two defect triangles, with the exception of segment 7, which contains only one. This segmentation strategy reflects both the topological structure of the Kagome lattice and the constraints of the hardware-efficient ansatz, allowing localized parameter optimization within manageable sub-circuit sizes.
A key objective of the local VQE phase is to determine a precise value of $J^{\prime}$ on selected bonds of each defect triangle. For every lattice segment, the value of $J^{\prime}$ is individually calibrated to ensure that the computed local ground-state energy matches the exact ground-state energy obtained via exact diagonalization. As shown in Table-2, the calibrated values of $J^{\prime}$ remain remarkably consistent across all subregions, clustering around $J^{\prime}\approx 1.95$. This empirical regularity supports the scalability of our Hamiltonian engineering approach to larger systems. At the conclusion of the local VQE phase, we obtain a baseline estimate of the ground-state energy gs_energy as $-161.98$ (see Table-2). In the subsequent phase, the sub-circuits that contributed to this value are recombined into a complete ansatz for global VQE, with the goal of further lowering the energy.
@@ IV.2 Global VQE: Quantum Computing
Local VQE produces pre-optimized ansatz segments for each subregion of the Kagome lattice. To recover the full 103-qubit circuit, these segments are concatenated such that all original gate and qubit dependencies of the ansatz are preserved, effectively reconstructing the near original circuit. However, this naïve concatenation does not guarantee global optimality because the junction qubits, shared between two segments, inherit independently optimized $R_{y}(\theta)$ rotations from each side. To address the potential sub-optimality introduced at these junctions, we insert an additional trainable $R_{y}(\theta)$ gate on each junction qubit (see Figure-3), and tune these parameters via a global VQE.
@@ IV.2.2 Objective Function
Concatenating seven local ansatz segments introduces six free parameters corresponding to the newly inserted $R_{y}(\theta)$ gates on the junction qubits (e.g., qubits 17, 33, 48, 63, 73, 92). The global VQE runs the reconstructed ansatz on IBM quantum processors, with the objective of minimizing $\text{Tr}(\rho H_{\text{pert}})$, where $H_{\text{pert}}$ comes from eq(III.1) and $\rho$ is the quantum state prepared by the ansatz circuit. The original Hamiltonian $H$ and $H_{\text{pert}}$ include over 507 Pauli product terms and each term accumulates noise when ansatz is readout on quantum processor, thereby causing unwanted fluctuations during global VQE. To address this problem we adopt a noise-aware simplification by retaining only those terms involving the junction qubits and their immediate neighbors, this reduced Hamiltonian is denoted $H_{\text{SEL}}$.
$H_{SEL}=J\sum_{(i,j)\in\mathcal{S}}X_{i}X_{j}+Y_{i}Y_{j}+Z_{i}Z_{j}$
@@
While this truncation reduces the expressiveness of the cost function, it significantly lowers the impact of noise. Owing to the entanglement path through the 102 CNOT gates of the hardware-efficient ansatz, all qubits remain technically interdependent, and thus parameter updates at a junction qubit gate $R_{y}(\theta)$ can, in principle, affect the quantum state of all the dependent qubits, as highlighted in the breakdown of ansatz circuit provided in Figure-3.
@@ IV.3 Noise Mitigation with Operator Decoherence Renormalization (ODR)
To implement this correction, we consider a parameter regime in which the VQE ansatz circuit becomes efficiently classically simulable, typically by reducing to a Clifford circuit. Importantly, the Clifford version of the ansatz arises solely due to specific parameter choices; the underlying circuit structure (i.e., the gates and their order) remains unchanged. We refer to this special configuration as the reference ansatz.
In our setting, the Clifford version of the reference ansatz prepares a static dimer cover state $|\phi\rangle$. On ideal (noiseless) quantum computer the circuit yields the expectation value $\langle\phi|H|\phi\rangle=-147$. On the other hand, when the reference ansatz circuit is executed on the quantum device, it produces a noisy (mixed) state $\sigma$, and the expectation $\operatorname{Tr}(\sigma H)$. In contrast, the original ansatz of the VQE algorithm typically operates in a non-Clifford regime, as it explores the full parameter space to minimize energy. Nevertheless, once the VQE optimization is complete, the ODR technique can be applied post hoc to mitigate the noise in the final gs_energy estimate.
@@
$\frac{\text{noise-mitigated}\operatorname{Tr}(\rho H_{\text{pert}})}{\operatorname{Tr}(\rho H_{\text{pert}})}=\frac{\langle{\phi}|H|\phi\rangle}{\operatorname{Tr}(\sigma H)}$
@@ V Experimental Results
Table-5 presents the optimal parameter sets obtained from the global VQE runs on ibm_aachen, ibm_fez, and ibm_torino. These optimal parameters appear to be significantly influenced by processor-specific factors, including calibration quality and error rates. Nevertheless, there is no fundamental limitation preventing cross-testing of these parameter sets across different processors, which can help expand the search space for identifying the lowest gs_energy estimate. Therefore, treating Table-5 as a general candidate pool, we proceed to compute the noise-mitigated $\textnormal{Tr}(\rho H_{pert})$ by trial assignment of each parameter set to the blue-shaded gates in Figure-3, across all processors.
The values reported in Table-3 were obtained by selecting, for each processor, the parameter set from Table-5 that yields the minimum noise-mitigated $\textnormal{Tr}(\rho H_{pert})$. To ensure consistency, we aim to maximize the ratio $\frac{\textnormal{Tr}(\rho H_{pert})}{\textnormal{Tr}(\sigma H)}$ within the same Qiskit job. This strategy avoids device-level behavioral and calibration fluctuations that can occur when measurements are spread across multiple Qiskit estimator jobs. The numerical values in Table-3 are based on at least three independent repetitions of the same configuration (circuit, parameter set, and estimator settings). See Figure-3(d) for details on the estimator settings. Table-3 summarizes energy estimates from five different quantum processors, fall between –165.91 and –172.32.
For comparison, classical simulations using Qiskit’s Aer backend with the Matrix Product State (MPS) method and circuit cutting [36] yield an estimate –167.95, closely matching the quantum version of the results. This validates the accuracy of our noise-mitigated global VQE. To benchmark our results, we introduce heuristic energy bounds. The 103-site system contains 60 triangles, each contributing –3 in the ideal dimerized configuration, giving a theoretical lower bound of –180. However, defect triangles systematically raise the energy. For example, a single defect triangle on the smallest sub-lattice (two-unit cells as in Figure-7(d)) elevates gs_energy by 0.86 (Table-2(d)). With 13 such defects, and each contributing $\approx$+0.86  the effective upper bound becomes –168.82. This value aligns closely with our best estimates from both quantum and classical simulations, reinforcing the accuracy of our Hamiltonian-engineered VQE strategy.
Table 3: The unmitigated (Actual) and noise-mitigated (Noise mitigated) gs_energy using ODR eq(2), for five different ibm quantum processors
Processor Actual $\textnormal{Tr}(\rho H_{pert})$ (units of $J$) Ref. ODR $\textnormal{Tr}(\sigma H)$ (units of $J$) Noise mitigated $\textnormal{Tr}(\rho H_{pert})$ (units of $J$)
@@ V.2 Comparison of Classically-optimized and Quantum-optimized Parameters
To ensure a fair comparison, both parameter sets were independently evaluated on three quantum processors, ibm_torino, ibm_marrakesh, and ibm_fez, by plugging them into the full 103-qubit ansatz and measuring the ground-state energy without applying any noise mitigation (e.g., ODR). The unmitigated energy values obtained from $\mathbf{p}_{1}$ were $-121.47\pm 1.05$, $-105.58\pm 0.72$, and $-115.77\pm 0.46$, respectively. In contrast, the energy values using $\mathbf{p}_{2}$ were consistently lower: $-122.81\pm 0.53$, $-106.63\pm 0.33$, and $-117.32\pm 0.22$. Although the numerical differences are moderate, these results suggest that parameters learned from quantum VQE directly on hardware not only reflect the underlying circuit structure but also adaptively account for calibration specifics and coherent noise characteristics [32], potentially navigating toward lower-energy configurations. This finding supports the view that quantum-optimized parameters can outperform those obtained from classical approximations, particularly when hardware-specific factors significantly influence algorithm convergence.
Table 4: Comparison of unmitigated ground-state energies using parameters from classical VQE ($\mathbf{p}_{1}$) and global (quantum) VQE ($\mathbf{p}_{2}$) on different IBM quantum processors.
Processor Energy with $\mathbf{p}_{1}$ (units of $J$) Energy with $\mathbf{p}_{2}$ (units of $J$) Average Energy Decrease (units of $J$)
@@
An important objective of this study is to demonstrate utility-scale quantum computation. We have already shown that the parameter set $\texttt{p}_{2}$, obtained via quantum execution, slightly outperforms the classical counterpart $\texttt{p}_{1}$ in one-to-one comparisons on specific processors. Figure-5 presents a more holistic, statistically informed comparison of global VQE parameter performance on quantum processors with that of $\texttt{p}_{1}$ on classical computer. The figure contrasts the quantum gs_energy estimates, averaged over all parameter sets listed in Table-5 and all five processors, with the classical baseline estimate using $\texttt{p}_{1}$ (i.e., $-167.95$).
@@
As quantum hardware continues to improve in gate fidelity and qubit coherence times, quantum validation of classically approximated quantum algorithms opens an exciting avenue for future research. It provides a concrete benchmark for assessing the reliability of classical simulations, especially as quantum processors scale to larger qubit counts with increasing gate fidelity. Our work presents an initial step toward this promising direction and encourages further exploration of its potential.
@@ VI Benchmarking Against the Thermodynamic Ground State Energy
In this work, we present what is, to our knowledge, the first experimental attempt to estimate the ground-state energy of a 103-site KAFH lattice on superconducting quantum hardware. By employing a Hamiltonian engineering strategy, we modulate bond strengths to favor valence bond configurations and optimize energy estimation using a hardware-efficient VQE ansatz. On IBM quantum processors, our estimate for the ground-state energy per site of a 103-site KAFH lattice requires conversion to the unscaled version of gs_energy becomes $-172.4J/4=-52.1J$ translating into per-site gs_energy $=-521.J/103=\texttt{-0.417}J$. This result lies above the thermodynamic-limit energy, which has been determined via large-scale DMRG and tensor network methods to lie between $-0.436$ and $-0.438J$ [13, 12].
@@
A substantial part of this discrepancy arises from finite-size and boundary effects, which are well known to raise energy estimates in systems with open boundary conditions. In our lattice geometry, 25 out of 125 sites lie on the boundary and possess only two nearest neighbors, as opposed to the four neighbors found in the bulk. This reduced local coordination diminishes quantum fluctuations and suppresses dimer resonance, leading to locally elevated bond energies and an increased average energy per site.
VI.1 Open Boundary Correction
To account for these edge effects, we apply an open-boundary correction (OBC) model, following established procedures from prior finite-size studies [28, 14]. We decompose the total energy of 125-site lattice into contributions from bulk and edge sites. The bulk sites have all four neighboring sites while edge sites lie on the boundary and have strictly fewer neighbors (usually 2). Following the approach used in prior DMRG and tensor network studies [7, 12, 44], we express energy as weighted sum of its bulk sites’ and edge sites’ energies as:
$E_{\text{total}}=\bar{E}_{\text{bulk}}N_{\text{bulk}}+\bar{E}_{\text{edge}}N_{\text{edge}}$ (3)
Where $N_{\text{bulk}}$ and $\bar{E}_{\text{bulk}}$ represent number and per-site energy of bulk sites while $N_{\text{edge}}$ and $\bar{E}_{\text{edge}}$ represent number and per-site energy of the edge sites. The corrected gs_energy, defined as $\bar{E}_{\text{bulk}}$, is given by following expression:
$\bar{E}_{\text{bulk}}=\frac{E_{\text{total}}-N_{\text{edge}}\cdot\bar{E}_{\text{edge}}}{N_{\text{bulk}}}$ (4)
Note that $\bar{E}_{\text{bulk}}$ and $\bar{E}_{\text{edge}}$ are obtained from the unscaled version of gs_energy divided by the number of sites. To use the eq(4), revert gs_energy to the un-scaled version and estimate the average edge-site energy $\bar{E}_{\text{edge}}$ using small lattice benchmarks, where exact diagonalization is possible for given $E_{\text{total}}$. Thus, for a 12-site Kagome unit cell, $E_{\text{total}}$ = $-18J/4=-4.5J$. Likewise, for the 19-site two-unit-cell system, we have $E_{\text{total}}=-29.14J/4=-7.285J$, and for the 23-site case (one site less than three full unit cells), $E_{\text{total}}=-36.43J/4=-9.1075J$. For each of these cases, eq(4) yields the corresponding edge-site energy values: $\bar{E}_{\text{edge}}=-0.3114J$, $-0.3076J$, and $-0.3296J$, respectively. Therefore, we estimate that $\bar{E}_{\text{edge}}$ lies within the range:
$\bar{E}_{\text{edge}}\in[-0.3076J,\,-0.3296J]$
For 125-site lattice, we calculate $\bar{E}_{\text{bulk}}$ minimum and maximum per-site energies. For a 125-site lattice our gs_energy becomes $-172.5-36=-208.4J$, which gives $E_{\text{total}}=52.1J$. The Figure-2 shows that $N_{\text{edge}}=25$ , $N_{\text{bulk}}=100$. For the range $[-0.3076J,-0.3296J]$, eq(2) implies
$\bar{E}_{\text{bulk}}(\text{max})=$ $-0.4386J$
$\bar{E}_{\text{bulk}}(\text{min})=$ $-0.4441J$
The bulk energy values reflect the underlying quantum correlations more accurately than the raw per-site average and closely match the one obtained in the thermodynamic limit $-0.4386J$ [13, 43],